# Agentic LMs: Hunting Down Test Smells

### v2

@@ 1 Introduction
Test smells are design flaws in test code that compromise reliability and hinder maintenance [1]. They can reduce the effectiveness of test suites. These issues are prevalent in both open-source and industrial settings [2]. Common test smells include undocumented assertions (Assertion Roulette), complex conditional logic (Conditional Test Logic), duplicated assertions (Duplicate Assert), and hard-coded values (Magic Number[2].
@@
Preliminary results demonstrate promising generalization across programming languages – specifically Python, Golang, and JavaScript – using the same setup. Beyond automating the detection and refactoring of test smells, our approach also enables developers to interact with models to better understand the rationale behind suggested changes. All artifacts are publicly available online. [^1] https://zenodo.org/records/16740114
@@ 3 Methodology
To remove a test smell, Agent3 utilizes the prompt outlined next. During the second or third iteration, it incorporates feedback from Agent4. You are a coding assistant specializing in test code analysis and refactoring, with many years of experience.
@@ 4 Evaluation
Table I summarizes the detailed performance (pass@1) of the single-agent and multi-agent approaches for detecting and refactoring test smells.
@@ 4.1 Detecting Test Smells
We review Agents1,2 outputs to assess the validity of their test smell explanations. Next, Agent1 uses Phi-4-14B to identify the Exception Handling test smell. The other models also detect it. “The test method testStackBlowOut includes a try-catch block to handle exceptions manually, which is indicative of the Exception Handling test smell. The code attempts to catch an IllegalArgumentException, rather than using testing framework features like assertThrows to verify that the exception is thrown as expected.” Llama-3.2-3B achieved the highest detection rate, identifying all test smell instances. Gemma-2-9B and Phi-4-14B missed a few but provided more detailed and accurate explanations. These results demonstrate that smaller models can reliably detect test smells when given concise natural language definitions. Unlike static-analysis tools, models take a semantic approach, recognizing frameworks like Mockito or JUnit without needing exhaustive method lists.
@@ 4.2 Refactoring Test Smells
The optimal configuration for Llama-3.2-3B and Gemma-2-9B involved using four-agents, successfully refactoring 22% and 40.7% out of 150 instances in one attempt, respectively. DeepSeek-R1-14B achieved its best performance with a single-agent, correctly refactoring 39.3% instances. In contrast, Phi-4-14B performed best with two-agents, successfully refactoring 55.3% instances, demonstrating its effectiveness in a multi-agent setup. Processing 150 smells using four-agents took ~1h (Llama-3.2-3B, Gemma-2-9B), ~5h (Phi-4-14B), and ~8h (DeepSeek-R1-14B).
@@
We can also correctly apply a refactoring to remove Duplicate Assert test smells using multi-agents with Phi-4-14B. However, using a single-agent with Phi-4-14B applies a refactoring strategy that differs from the one proposed for this smell. Specifically, it incorporates a unique message for each assertion, effectively eliminating both this test smell and the Assertion Roulette test smell.
@@
For the Assertion Roulette test smell, Phi-4-14B achieved the best result, correctly detecting and refactoring 93.3% of instances with a single-agent. In contrast, all models struggled with Conditional Test Logic, where we evaluated only one of several possible refactoring strategies. For Duplicate Assert, the highest success rate (70%) was obtained by Phi-4-14B using two-agents. Phi-4-14B also performed best on Exception Handling, correctly refactoring 43.3% of cases with a single-agent. The best results for Magic Number came from DeepSeek-R1-14B with two agents, achieving 90% accuracy. Overall, two or four-agent setups led to better outcomes in three of the five test smells. However, for Assertion Roulette and Exception Handling, a single-agent was sufficient. Notably, there are 23 unique instances successfully refactored by at least one model – Llama-3.2-3B (6), Gemma-2-9B (14), or DeepSeek-R1-14B (6) – that Phi-4-14B with two-agents failed to refactor. By combining strengths across models, we achieve correct refactorings for 70.6% instances (pass@1).
@@ 4.3 Failure Types
The ablation study revealed distinct failure patterns across agent configurations. In the single-agent setup, issues frequently included incomplete fixes, superficial code changes that failed to resolve the smell, and the introduction of invalid or non-existent constructs—often stemming from incorrect syntax or reliance on frameworks other than JUnit-5. In Magic Number, the model sometimes extracted only a subset of constants or failed to initialize new variables. Refactorings for Conditional Test Logic and Duplicate Assert occasionally introduced subtle logic changes. In Exception Handling, some transformations mistakenly removed critical code alongside the try block.
Introducing additional agents reduced some errors but brought new challenges, including increased code divergence, missed detections that the single agent previously succeeded on, and frequent disagreements—particularly with Agent4 rejecting correct refactorings proposed by Agent3. Moreover, if the multi-agent setup had used JUnit-5 instead of JUnit-4 and correctly positioned the message parameter, its performance in Assertion Roulette would have matched that of the single-agent configuration.
@@ 4.4 Feedback Loop
The feedback loop was triggered when Agents2,4 disagreed with Agents1,3, respectively (see Figure 1). For detecting test smells, Gemma-2-9B and Phi-4-14B utilized the feedback loop in 1 and 9 test smell instances, respectively. In 2 instances, Phi-4-14B was able to correctly identify a test smell due to the feedback loop. Llama-3.2-3B did not require the feedback loop for detecting any test smells. DeepSeek-R1-14B successfully detected 3 instances of the Duplicate Assert test smell following a feedback loop. We employed a three-iteration feedback loop, which was sufficient for Agents1,2 to reach a consensus on detecting the majority of test smells. For Agents3,4 to reach a consensus on refactoring test smells, only 7.3% of the cases required more than three iterations. By allowing additional iterations, we achieved consensus in all cases by the eighth iteration, except for a single instance of the Conditional Test Logic test smell.
@@ 4.5 Pull Requests
To assess the acceptance of the refactorings generated by Phi-4-14B, which achieved the best performance in our study, we submitted 10 pull requests, ensuring at least one for each test smell type across five different open-source projects. As of this writing, six of these pull requests have been reviewed and merged by human maintainers, who deemed them valuable: two in janusgraph (Assertion Roulette and Exception Handling), two in lettuce (Magic Number and Duplicate Assert), and one each in opengrok and jenkins (both addressing Duplicate Assert). Two pull request were rejected because contributors noted that the additional message in each assertion did not improve their code. Two pull requests remain open.
@@ 4.8 Other Languages
We assess whether the same setup can detect and refactor test smells across different programming languages. Using a single-agent configuration, Phi-4-14B successfully refactored (pass@1) 24/30, 27/30, and 29/30 test smells in Golang, JavaScript, and Python, respectively. These instances contained Assertion Roulette, Duplicate Assert, and Magic Number smells in real-world projects such as kubernetes, react-beautiful-dnd, and django. The model achieved performance comparable to that observed with Java test cases. In the JavaScript setting, it occasionally generated Java code; however, when explicitly instructed to produce JavaScript test code, it responded correctly. JUnit-5 compliance can also be excluded from the prompts. These preliminary results suggest that the proposed approach is promising for detecting and refactoring test smells across a variety of programming languages.
@@ 4.10 Larger Test Cases
As previously noted, 89% of the test cases in the 11 open-source projects analyzed contain at most 30 lines of code (LOC). To further evaluate Phi-4-14B, which demonstrated the best performance in our study, we selected 30 test smells (five per smell type) exceeding 30 LOC from these 11 open-source projects. We employed the same four-agent setup to detect and refactor these test smells. Although the strategy successfully identified 86% of the test smells, it showed a decline in performance compared to previous results. The effectiveness in accurately refactoring test smells also diminished, with a pass@1 rate of 28%.
@@ 6 Conclusions
We evaluated Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B for the automated detection and refactoring of test smells using agentic workflows with one, two, and four agents. The evaluation covered 150 instances of 5 common test smell types from real-world projects. We assessed each model’s ability to perform automated refactorings. Phi-4-14B, Gemma-2-9B, and Llama-3.2-3B successfully detected nearly all test smell instances. Phi-4-14B delivered the best refactoring performance, reaching a pass@5 of 75.3%, and performed within 5% of proprietary models like o3, Claude-4-Sonnet, and Gemini-2.5-Pro using a single-agent. We submitted 10 pull requests with code refactored by Phi-4-14B. Six were merged, two was rejected, and two remain open – demonstrating the practical utility of our approach in live development environments. Multi-agent workflows outperformed single-agent setups in three out of five test smell types, showing their potential to enhance code quality with minimal developer effort.

### v3

@@ 1 Introduction
Test smells are design flaws in test code that compromise reliability and hinder maintenance [1]. They can reduce the effectiveness of test suites. These issues are prevalent in both open-source and industrial settings [2]. Common test smells include undocumented assertions (Assertion Roulette), complex conditional logic (Conditional Test Logic), duplicated assertions (Duplicate Assert), and hard-coded values (Magic Number) [2].
@@
Preliminary results demonstrate promising generalization across programming languages – specifically Python, Golang, and JavaScript – using the same setup. Beyond automating the detection and refactoring of test smells, our approach also enables developers to interact with models to better understand the rationale behind suggested changes. All artifacts are publicly available online. [^1] https://zenodo.org/records/17285750
@@ 3 Methodology
To remove a test smell, Agent3 utilizes the prompt outlined next. During the second or third iteration, it incorporates feedback from Agent4.
You are a coding assistant specializing in test code analysis and refactoring, with many years of experience.
@@ 4 Evaluation
Table I summarizes the detailed performance (pass@1) of the single-agent and multi-agent approaches for detecting and refactoring test smells. In our setting, pass@1 measures whether the model’s first attempt correctly detects or refactors a test smell, with correctness independently verified by two authors and cross-checked by a third.
@@ 4.1 Detecting Test Smells
We review Agents1,2 outputs to assess the validity of their test smell explanations. Next, Agent1 uses Phi-4-14B to identify the Exception Handling test smell. The other models also detect it.
“The test method testStackBlowOut includes a try-catch block to handle exceptions manually, which is indicative of the Exception Handling test smell. The code attempts to catch an IllegalArgumentException, rather than using testing framework features like assertThrows to verify that the exception is thrown as expected.”
Llama-3.2-3B achieved the highest detection rate, identifying all test smell instances. Gemma-2-9B and Phi-4-14B missed a few but provided more detailed and accurate explanations. These results demonstrate that smaller models can reliably detect test smells when given concise natural language definitions. Unlike static-analysis tools, models take a semantic approach, recognizing frameworks like Mockito or JUnit without needing exhaustive method lists.
@@ 4.2 Refactoring Test Smells
The optimal configuration for Llama-3.2-3B and Gemma-2-9B involved using four-agents, successfully refactoring 22% and 40.7% out of 150 instances in one attempt, respectively. DeepSeek-R1-14B achieved its best performance with a single-agent, correctly refactoring 39.3% instances. Phi-4-14B performed best with two-agents, successfully refactoring 55.3% instances, demonstrating its effectiveness in a multi-agent setup. Processing 150 smells using four-agents took ~1h (Llama-3.2-3B, Gemma-2-9B), ~5h (Phi-4-14B), and ~8h (DeepSeek-R1-14B).
@@
We can also correctly apply a refactoring to remove Duplicate Assert using multi-agents with Phi-4-14B. However, using a single-agent with Phi-4-14B applies a refactoring strategy that differs from the one proposed for this smell. It incorporates a unique message for each assertion, effectively eliminating both this test smell and the Assertion Roulette test smell.
@@
For Assertion Roulette, Phi-4-14B achieved the best result, correctly detecting and refactoring 93.3% of instances with a single-agent. In contrast, all models struggled with Conditional Test Logic, where we evaluated only one of several possible refactoring strategies. For Duplicate Assert, the highest success rate (70%) was obtained by Phi-4-14B using two-agents. Phi-4-14B also performed best on Exception Handling, correctly refactoring 43.3% of cases with a single-agent. The best results for Magic Number came from DeepSeek-R1-14B with two agents, achieving 90% accuracy. Overall, two or four-agent setups led to better outcomes in three of the five test smells. However, for Assertion Roulette and Exception Handling, a single-agent was sufficient. Notably, there are 23 unique instances successfully refactored by at least one model – Llama-3.2-3B (6), Gemma-2-9B (14), or DeepSeek-R1-14B (6) – that Phi-4-14B with two-agents failed to refactor. By combining strengths across models, we achieve correct refactorings for 70.6% instances (pass@1).
@@ 4.3 Failure Types
The ablation study revealed distinct failure patterns across agent configurations. Although adding more agents reduced certain errors, it also introduced new challenges, such as missed detections previously handled by the single-agent approach, incomplete fixes, and cases where Agent4 rejected correct refactorings proposed by Agent3. In the aggregated analysis, the multi-agent setup outperformed the single-agent configuration in three test smells, matched its performance in Exception Handling, and underperformed only in Assertion Roulette. The instance-level analysis with Phi-4-14B identified 12 cases in which the single-agent succeeded but the multi-agent failed. These included missed detections in Conditional Test Logic and Magic Number; misplaced or invalid refactorings in Assertion Roulette, where the model sometimes generated JUnit-4 instead of JUnit-5 messages; partial fixes in Magic Number, where only a subset of constants was extracted or new variables were not properly initialized; and issues in Exception Handling, where critical code was inadvertently removed along with the try block. In addition, one Duplicate Assert case introduced a behavioral change.
@@ 4.4 Feedback Loop
The feedback loop was triggered when Agents2,4 disagreed with Agents1,3, respectively (see Figure 1). For detecting test smells, Gemma-2-9B and Phi-4-14B utilized the feedback loop in 1 and 9 test smell instances, respectively. In 2 instances, Phi-4-14B was able to correctly identify a test smell due to the feedback loop. Llama-3.2-3B did not require the feedback loop for detecting any test smells. DeepSeek-R1-14B successfully detected 3 instances of the Duplicate Assert test smell following a feedback loop. We employed a three-iteration feedback loop, which was sufficient for Agents1,2 to reach a consensus on detecting the majority of test smells. For Agents3,4 to reach a consensus on refactoring test smells, only 7.3% of the cases required more than three iterations. By allowing additional iterations, we achieved consensus in all cases by the eighth iteration, except for a single instance of Conditional Test Logic.
@@ 4.5 Pull Requests
To assess the acceptance of the refactorings generated by Phi-4-14B, which achieved the best performance in our study, we submitted 10 pull requests, across five different open-source projects. As of this writing, six of these pull requests have been reviewed and merged by human maintainers, who deemed them valuable: two in janusgraph (Assertion Roulette and Exception Handling), two in lettuce (Magic Number and Duplicate Assert), and one each in opengrok and jenkins (both addressing Duplicate Assert). Two pull requests were rejected because contributors noted that the additional message in each assertion did not improve their code. Two pull requests remain open.
@@ 4.8 Other Languages
We assess whether the same setup can detect and refactor test smells across different programming languages. Using a single-agent configuration, Phi-4-14B successfully refactored (pass@1) 24/30, 25/30, and 29/30 test smells in Golang, JavaScript, and Python, respectively. These instances contained Assertion Roulette, Duplicate Assert, and Magic Number smells in real-world projects such as kubernetes, react-beautiful-dnd, and django. The model achieved performance comparable to that observed with Java test cases. In the JavaScript setting, it occasionally generated Java code; however, when explicitly instructed to produce JavaScript test code, it responded correctly. These preliminary results suggest that the proposed approach is promising for detecting and refactoring test smells across a variety of programming languages.
@@ 4.10 Larger Test Cases
As previously noted, 89% of the test cases in the 11 open-source projects analyzed contain at most 30 LOC. To further evaluate Phi-4-14B, which demonstrated the best performance in our study, we selected 50 test smells (ten per smell type) exceeding 30 LOC from these 11 open-source projects. We employed the same four-agent setup to detect and refactor these test smells. Although the strategy successfully identified 86% of the test smells, it showed a decline in performance compared to previous results. The effectiveness in accurately refactoring test smells also diminished, with a pass@1 rate of 28%.
4.11 Code Smells
As a feasibility study, we evaluated Phi-4-14B on 30 Java code smells drawn from real-world projects. By making minor adjustments to the prompts—for example, replacing references to “test” with “code”—Phi-4-14B successfully detected all instances and correctly refactored a subset of them. These results indicate that the approach is promising for the automated detection and refactoring of code smells. We leave a deeper investigation to future work.
@@ 6 Conclusions
We evaluated Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B for the automated detection and refactoring of test smells using agentic workflows with one, two, and four agents. The evaluation covered 150 instances of 5 common test smell types from real-world projects. We assessed each model’s ability to perform automated refactorings. Phi-4-14B, Gemma-2-9B, and Llama-3.2-3B successfully detected nearly all test smell instances. Phi-4-14B delivered the best refactoring performance, reaching a pass@5 of 75.3%, and performed within 5% of proprietary models like o3, Claude-4-Sonnet, and Gemini-2.5-Pro using a single-agent. We submitted 10 Phi-4-14B-refactored pull requests, and six were merged, showing practical utility. Multi-agent workflows outperformed single-agent setups in three out of five test smell types.
Acknowledgments
We thank the reviewers for their feedback. This work was partially supported by CNPq (403719/2024-0, 310313/2022-8, 404825/2023-0, 443393/2023-0, 312195/2021-4), FAPESQ-PB (268/2025), and FAPESB.